Error measurement of volume of a sphere is 1%.The error in calculation of surface area of the sphere will be what?

Grade:

1 Answers

Sanjeev Malik IIT BHU
12 years ago

vol= 4nr3/3

error can be in measuring radius only.

as power is 3 to r so errors is added 3 times.

so total error(max) can be 3*(error in r)

3*error=1%

error in r=1/3%

surface are = 4nr2=  so error is 2* error=  2*1/3%

=2/3%

 

please approve by clicking yes below if helpful